

********************************************************************
UARS WEEKLY STATUS REPORT - 05 June 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM has successfully gathered its normal solar, offset, and
occultation scan data without apparent problems.
We expect to place a new version of the SUSIM production software into
CDHF Engineering mode during the coming week.
***********
PEM
***********
PEM/AXIS was active from May 22 through May 23. It generated energy
deposition from inverted X-ray spectra for each of its 16 look directions.
GSFC has proposed an OBC operation plan for the activation of PEM boom
instruments during the daylight portions of every orbit. This OBC operation
plan is accepted by PEM. PEM has proposed a series of test procedures for
activation and is waiting response from GSFC. PEM is also waiting for GSFC
scheduling and implementation of the OBC plan and its operation test procedures.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
At the end of the last operating period for MLS, the MLS instrument
experienced a minor recurrence of the scan slips last observed several years
ago. These occurred during the last 3 periods of lowest bus voltage just prior
to normal instrument turn off.
Close examination of the Level 0 data indicates that
1. all scan slips took place during retrace, when the scan motor is known to
have lowest torque,
2. no additional wear is evident in the scan mechanism compared to data taken
several years ago, and
3. recovery from the anomalies was handled promptly and correctly by the
on-board software.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Gordon Shepherd is currently in Boulder, at HAO/NCAR, working with
Ray Roble on WINDII comparisons with the TIME-GCM.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 390 CD-ROM requests as of 04-June-1998.
There were no new orders received.
The problem encountered by the UK Meteorological Office at an intermediate
point has been corrected. The CDHF is currently up to date receiving the
UKMO data.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) PROD (1379R-1313C=66L)
61 days HAL (4.80) PROD (L1) (1868R-311C-2F=1555L)
31 days HAL (4.80) PROD (L3) (1868R-75C=1793L)
611 days SOL (4.61) PROD (L1) (2404R-1154C=1250L)
93 days SOL (4.61) PROD (L3) (2304R-169C=2135L)
1 day SUS (6.3) PROD (1R-1C=0L)
1 day HRDI(4.4) PROD (1R-1C=0L)
CDHF pending software upgrades: IDL 5.1, OpenIngres
***********
FOT
***********
REPORTING PERIOD
This report covers 23 May 1998 (Orbit 36588, GMT Day 143) through 29 May
1998 (Orbit 36692, GMT Day 149).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 143/0754 36593 Axis Off*
WINDII 143/0754 36593 Off*
HRDI 143/2333 36603 Low Pwr Mode
WINDII 144/2243 36617 On
WINDII 145/1316 36626 Off
* TMON 01 activation
The beta angle reached a maximum angle of 78.0 degrees on 26 May 1998 (DOY
146 and is now decreasing toward a minimum angle of 0.0 degrees on 15 June
1998 (DOY 166).
On Board Computer (OBC) Telemetry Monitor (TMON) 01 tripped in Orbit 36593
on 23 May 1998 (143/0745) due to low voltage (23.36 volts). This resulted
in Relative Time Sequence (RTS) 00 activation which powered OFF an energy
balance deficit and TMON 00 trip which commanded off WINDII, HRDI, and PEM.
Power balance was restored after TMON 01 tripped. HRDI returned to low
power in orbit 36603 (143/2333). WINDII was turned back on in Orbit 36617
(144/2243), but had to be turned back off in Orbit 36626 (145/1316) due to
power balance deficit. The SSPP instruments and MLS remained ON throughout
the high solar beta range above 73 degrees. The spacecraft had increased
thermostatic heater activity at the high solar beta angle, and coupled with
near minimum solar array output, was unable to support the scheduled
instrument loads.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.6 V end-of-night (EON) load bus voltage.
Battery temperatures are increasing (due to high beta angle and low depth-
of-discharge) with a temperature delta between Battery 2 and Battery 3 of
1.0 degrees C. Additional battery performance data for each battery on 23
May 1998 and 30 May 1998 are provided in Appendix D.
The clock error ranged from -10.6 to +11.7 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 012
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 04
% HRDI 02
% PEM 01
% WINDII 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 03
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/118 - 2287.497007
Transponder B: 98/120 - 2287.498907
GROUND SYSTEM OPERATION
Two (2) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/148 show the observatory data loss to be
20 hours, 43 minutes, 29.652 seconds (no change since last report period).
This is a 0.0354 percent data loss which equals a 99.9646 percent data
capture for the mission.
The FOT made a few more modifications to the OBCSTF resident POCC database
this week. These changes were necessary to accommodate the proposed HRDI
"day-only" operations scenario.
The FOT made a few additional modifications to the OBCSTF resident POCC
database this week. These changes were necessary to accommodate the
proposed HRDI "day-only" operations scenario. Due to unavailability of the
MSOCC software coordinator this week, testing of the latest version of
database 67 will most likely be delayed until next week.
Testing the command side of the I/P Net remains on hold. Testing will
begin when MSOCC provides a command loop interface from the POCC through
the I/P Net to the UARS Simulator.
OTHER SIGNIFICANT EVENTS
Continued testing of the upgrade to the OBC Power Monitor (PMON) flight
software. PMON software modifications and testing continued this week.
Multiple simulations on the new AP software including both the new database
version 67 and the old version 65. Highlights of the testing involved
integration and validation of HRDI commands into the new power-orbit action
plan. Minor changes to other instrument RTSUs were required.
Work continues on the UARS orbit adjust planning scheduled for 05 August
1998. Additional options for an Orbit Adjust are being planned in the
event that the B/D thrusters can not be safely used.
UARS continues to assiste TOMS in reviewing the TOMS/EP Transmitter A
failure data.
FOT continues to support CNMOS in their efforts to make necessary software
changes (MSOCC, CMS, DCF, etc.) to support the year 2000 rollover. The FOT
requested a set of "test" planning aids from FDF to facilitate Year 2000
ground system testing. FDF informed the FOT that they are currently in the
process of generating planning aids that span the Year 2000 crossover.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 15 June 1998.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 05 August 1998.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 29 May 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
36594 143/0937 98-176 PEM Inhibit Mission Unique Bit
36600 143/1913 98-177 HRDI Power ON to Low Power
36603 143/2333 98-178 HRDI Pulse Bus Off
36617 144/2243 98-179 OBC Clk Adj from 41.2 to 20.6 msec/day
36617 144/2243 98-180 WINDII Power ON
36625 145/1122 98-181 PSU Comp Htr 18 A/B OFF
36626 145/1316 98-182 WINDII WINDII TURN OFF
36638 146/0805 98-183 OBC Clk Adj from 20.6 to 41.2 msec/day
36689 149/1754 98-184 OBC Clk Adj fr 41.6 to 20.6 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 29 May 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIPBVOLTAGE Y-LO WINDII Pulse Bus Voltage 36592,36625
HRPSEQBIMON Y-HI HRDI Q/B Current Monitor 36592,36623-25,28
CDUNRG28 Y-LO Unregulated +28 Voltage 36592-3,36624,25
PWLBV Y-LO Load Bus Voltage 36593
WIQBVOLTAGE Y-LO WINDII Quiet Bus Voltage 36624
PWBAT3V Y-LO Battery 3 Voltage 36593
PWBAT2V Y-LO Battery 2 Voltage 36593
OBSVLBEON Y-LO Last EON Load Bus Voltage 36593-4,36626
OBSBATVEON2 Y-LO Battery 2 EON Voltage 36593-4
OBSBATVEON3 Y-LO Battery 3 EON Voltage 36593-4
HRIFESURVTMP Y-HI IFE Survival Heater Temp 36594
HRRADTEMP Y-HI IFE Radiator Temp 36594
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 29 May 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
36617 144/22:48 (24 May 98) from 00.0 to 20.6 msec/day
36638 146/20:44 (26 May 98) from 20.0 to 41.2 msec/day
36689 149/17:35 (29 May 98) from 41.2 to 20.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 29 May 1998
23 May 1998 - GMT Day 143 Beta = 73.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.94/-4.64 +0.0/ +0.0 18.2 0.000 0.0
2 +42.0/ +0.0 +3.87/+2.10 +17.1/-16.1 23.5 1.043 21.8
3 +67.6/ -5.6 +3.06/+1.34 +16.8/-16.2 23.4 1.045 21.3
29 May 1998 - GMT Day 149 Beta = 71.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.4 0.000 0.0
2 +44.8/ -11.2 +3.42/+2.05 +18.0/-15.2 26.9 1.039 18.3
3 +67.5/ -5.5 +2.50/+1.34 +17.3/-14.4 26.7 1.038 18.1
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 29 May 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-025 FOT 36592 143/0616 19919 Incomplete Config. TDRS
98-021 FOT 36603 144/0107 14170 Generic Late Acquisition
98-026 FOT 36680 149/0315 TAC #4 Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-025 FOT 36592 143:06:16 19919 WHITE SANDS CONF.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-021 Generic Late Acq for May 1998 (Open until 31 May 1998).